# Are AI-Generated Fixes Secure? Analyzing LLM and Agent Patches on SWE-bench

### v2

@@ IV Threats to Validity
Our security analysis relies on static analysis tools, which are known to produce false positives and vary in coverage. To mitigate this threat, we used three tools (Semgrep, Bandit, and CodeQL) and applied a majority voting technique to reduce false positives and increase confidence in the identified vulnerabilities. This approach deliberately trades off recall for precision; while we detect fewer vulnerabilities than we might using a single tool or a union-based strategy, we gain higher confidence that the findings represent true security issues. Nevertheless, we acknowledge that some false positives or missed vulnerabilities may remain.
@@
Moreover, our evaluation focused on one high-performing LLM and the top three publicly available agentic frameworks at the time. While there are many other models and frameworks in the space, our selection captures a broad and representative sample of state-of-the-art systems. Further, while our LLM-based approach generated responses for both the train and test sets (a process that took well over a week), we acknowledge the limitation of our agentic study in terms of data scale and have adjusted our conclusions accordingly. For example, AutoCodeRover reports a cost of \$0.70 per instance to achieve their test set results, meaning it would cost over \$13,000 to generate patches for the full 19,000 instances of training set, which is beyond our available resources.
@@
Lastly, manual annotations may be prone to human error. To improve reliability, all annotations were independently performed by two reviewers, whom ensured a strong inter-rater agreement (Cohen’s Kappa $>$ 0.9) through discussions.
V Related Work
@@
Security of LLM-generated Code. Given the growing adoption of LLMs by developers [60, 61, 62], recent studies have begun to examine various aspects of LLM-generated code [63, 64, 23, 35]. While many of these studies have focused on functionality, others have examined the security implications of LLM-generated code [13, 14, 51, 19, 18, 17]. Pearce et al. [13] conducted one of the first assessments on the security of LLM-generated code, showing that even state-of-the-art models can produce vulnerable code in security-sensitive scenarios. Since then, many studies have confirmed these findings by analyzing LLM-generated code across different tasks, models, and user interaction settings [65, 9, 66, 12].
@@
Most prior work has focused on small code snippets or controlled experimental settings [17, 18, 13, 19]. These settings often lack the full project context and complexity inherent in real-world software maintenance tasks, limiting their ability to assess how vulnerabilities emerge in practice. However, a recent study by Fu et al. [14] analyzed Copilot-generated code snippets in GitHub projects and identified several security weaknesses. It is worth mentioning that because these completions are reviewed and integrated by developers, the model is not fully in control of the final output. Our work complements and extends this work by analyzing LLM-generated code in real-world GitHub issues, where models are given autonomy to modify any relevant files. To our knowledge, this is the first large-scale empirical study to link vulnerabilities introduced by LLMs and agentic frameworks to a broad set of contextual factors, including issue descriptions, code-level attributes, and project structure, in realistic software development settings where models are tasked with resolving issues.
@@
Abnormal Commit Detection. Past studies have explored various techniques to identify and characterize the risk of potential changes in software systems, particularly focusing on GitHub commits [48, 46, 67, 49]. This line of research will generally aim to identify outlier commits, also known as risky or abnormal commits, through assessing various commit attributes. For instance, Gonzalez et al. introduced Anomalicious [49], an approach to identify malicious commits. By analyzing the commits and identifying the characteristics of a “normal” developer commit, they made an attempt at detecting malicious code contributions.
@@
VI Conclusions
This study presents the first large-scale security evaluation of LLM and agent-generated patches for real-world issue resolution tasks. We find that standalone LLMs can introduce up to 9× more new vulnerabilities than developers when addressing issues. Notably, some of the most frequent vulnerabilities introduced by LLMs (e.g., CWE-95, CWE-327) differ significantly from those found in developer-written code (e.g., CWE-732, CWE-377), indicating that LLMs are not merely mimicking human mistakes but also introducing distinct patterns of failure. Further, our analysis of agentic frameworks shows that increasing LLM autonomy can further amplify vulnerability risks, especially in tasks that are underspecified or require an understanding of the existing project context.
Importantly, our findings show that vulnerable outputs are not random. Rather, they are associated with particular code and issue-level characteristics, such as a higher number of modified files, longer generated code, and the absence of code snippets or certain information in the issue description. These patterns suggest that the riskiness of LLM or agent-generated patches can be assessed not only by detecting specific vulnerable lines of code but also by developing complementary risk assessment techniques that evaluate both the generated code and the broader context in which a patch is proposed (e.g., issue type or information completeness). Such systems can complement existing vulnerability detection techniques, often reliant on costly manual reviews, by narrowing the focus to potentially high-risk code contributions.

### v3

@@ 4. Threats to Validity
Our security analysis uses static analysis tools, which can produce false positives and vary in coverage. To mitigate this threat, we used three tools (Semgrep, Bandit, and CodeQL), maximizing coverage, and applied a majority voting technique to reduce false positives and increase confidence in the identified vulnerabilities. We further, performed a manual analysis on all the instances to remove all remaining false positives.
@@
Moreover, our evaluation focused on one high-performing LLM and the top three publicly available agentic frameworks at the time. While there are many other models and frameworks in the space, our selection captures a broad and representative sample of state-of-the-art systems. Additionally, while our LLM-based approach generated responses for both the train and test sets (a process that took well over a week), we acknowledge the limitation in terms of data scale and have adjusted our conclusions accordingly. For instance, AutoCodeRover reports a cost of \$0.70 per instance to achieve their test set results, meaning it would cost over \$13,000 to generate patches for the full 19,000 instances of training set, which is beyond our available resources. To keep our conclusions accurate, we therefore refrain from making strong statistical claims about the agentic systems, report their raw counts transparently, focus on qualitative analysis of the vulnerable instances they produce, and avoid direct quantitative comparisons with the standalone LLM.
We also acknowledge that our RQ3 analysis of issue types is limited by scale as manually annotating all 20,000 SWE-bench issues is infeasible, and many issues lack reliable labels. As a result, we do not claim that bug-fixing issues are inherently more prone to vulnerable patches; we only report that, within our manually annotated vulnerable instances, most affected issues are labeled as bugs.
@@
Lastly, manual issue annotations may be prone to human error. To improve reliability, all annotations were independently performed by two reviewers, whom ensured a strong inter-rater agreement (Cohen’s Kappa >> 0.9).
5. Discussion
5.1. Llama vs. Developers (RQ1)
The RQ1 results show that Llama introduces substantially more new vulnerabilities than developers on the same SWE-bench issues, even after we apply a conservative detection pipeline with majority voting and manual validation. Qualitatively, the vulnerabilities in LLM-generated code differ from developer-written patches. Both introduce issues such as command injection and path traversal, but LLM patches disproportionately exhibit categories like eval injection and insecure deserialization, and in some cases downgrade existing cryptographic mechanisms or include multiple vulnerabilities within a single patch. These patterns indicate that, in an APR setting, LLMs do more than repeat common developer mistakes; rather, they demonstrate distinct vulnerability patterns when asked to repair real-world bugs.
These findings indicate that Llama’s patches should be treated as an additional source of security risk. From an adoption standpoint, this indicates a clear trade-off: the LLM can accelerate development, but it also introduces a non-trivial probability of latent vulnerabilities, even in mature projects. This implies that (i) generated patches should not bypass existing code review and security processes, (ii) organizations deploying LLM-based APR tools should consider integrating security scanning and some form of manual inspection into workflows that accept LLM patches, and (iii) using these LLM fixes as a replacement for developer-written patches, without extra safeguards, is likely to increase both the frequency and diversity of vulnerabilities in codebases.
5.2. Agentic APR Frameworks (RQ2)
The RQ2 results show that agentic APR workflows built around LLMs do introduce security vulnerabilities, and that these vulnerabilities are not evenly distributed across frameworks or issues. In particular, OH’s patches display significantly different behavior than those of ACR and HC. A small subset of issues result in very large, multi-file diffs and multiple confirmed vulnerabilities produced by OH, whereas the other frameworks tend to modify fewer files and produce smaller patches.
This pattern points to a problem of highly autonomous APR frameworks. For an agent with broad permissions to edit the repository, run arbitrary commands, and generate new scripts for hundreds of iterations, any initial incorrect assumption regarding the codebase or the task can lead to long ineffective attempts to fix a problem. Over iterations, the patch may grow in size and scope, drifting away from the project’s original design and security expectations by modifying components outside the fault-localization region that the agent does not fully comprehend.
The comparatively small number of vulnerabilities observed in ACR and HC should not be interpreted as evidence that these systems are inherently secure, but it can suggest that tighter control over search space and modification scope can reduce the likelihood of vulnerabilities. ACR, particularly, operates with more constrained edit patterns and less extreme diff sizes than OH, which plausibly limits opportunities for large incorrect rewrites. At the same time, the vulnerabilities these frameworks introduce often fall into the same CWE categories e.g., command injection, eval injection, indicating that the core security issues of the underlying LLMs persist even when used in more restricted frameworks.
The results of RQ2 suggest that (i) giving agents unconstrained autonomy over large repositories can substantially increase the risk of vulnerabilities, especially via large, cross-file edits; (ii) frameworks that keep their edits more localized appear to have smaller vulnerability counts; and (iii) in all cases, the vulnerabilities introduced by agentic frameworks are distinct from those found in developer patches (similar to RQ1), emphasizing that these tools add a new layer of risk rather than merely repeating developers’ mistakes.
For practitioners considering agentic APR tools, these findings point to potentials for important design and deployment guidelines. Frameworks should incorporate explicit limits on edit scope e.g., caps on diff size, limits on the number of files that can be modified in a single attempt, etc. Additionally, given that a small number of issues can account for a disproportionate share of vulnerabilities, it is beneficial to detect high-risk repair tasks such as those that repeatedly fail tests or result in unreasonably large diff files, and route them to human review rather than allowing the agent to continue autonomously.
5.3. Code, Issue, and Project Characteristics of Vulnerable Instances (RQ3)
Our analysis of over 20,000 LLM-generated patches suggests that vulnerability-prone outputs are not random. Instead, they appear around a relatively consistent set of characteristics, especially at the code and issue level.
At the code level, we see a strong association between insecure patches and broader, more scattered edits. Vulnerable patches tend to touch more files and spread their changes across multiple locations, rather than making a self-contained change inside a single function or module. These are the situations where an LLM or APR agent must reason about cross-file dataflows and the interaction between different components, rather than a small local context. Our results indicate that such multi-file, dispersed edits are where current systems are most likely to introduce security weaknesses. In practice, if a repair task can be localized to a single file or a specific component, allowing an LLM to propose a patch in that narrow context appears substantially safer than giving it free rein over a large part of the repository. When tasks require multi-file edits, it may be preferable to decompose the work into smaller changes, each operating over a part of the code, rather than requesting one large cross-file patch. Additionally, we observe that in ACR, vulnerable patches are associated with significantly higher cyclomatic complexity. This is not a persistent signal across all systems, indicating that risk factors can depend on the internal workflow of the APR system. This suggests that for some APR frameworks, complexity-sensitive safeguards, e.g., requiring human review or additional checks when edits occur in highly complex functions, may be justified, even if the same signal is less informative elsewhere.
At the issue level, vulnerabilities are commonly found in bugs where the issue descriptions lack information such as clear statements about the observed and expected behavior or steps to reproduce the bug. This shows the importance of providing APR systems with complete and precise problem descriptions in order to minimize the risk of insecure patches. Interestingly, across both Llama and the frameworks, higher word counts in issue descriptions do not strongly distinguish vulnerable from non-vulnerable instances, indicating that the quality and relevance of the information can be a more significant factor than sheer verbosity. Long stack traces, or unrelated debugging conversations do not necessarily help the LLM, and may even divert its attention away from the important details.
At the project level, we do not observe strong correlations between vulnerability introduction and repository size, code complexity, maintainability, or contributor count. This may seem counterintuitive as one might expect that larger or more complex projects are inherently harder to repair securely. However, modern issue-tracking and modular design practices often constrain a given fix to a specific feature, module, or service, so many repairs remain local even in large systems. Conversely, a small codebase with tangled responsibilities and high cross-file coupling can force the model to reason across many interacting components even for a relatively simple problem.
Overall, the fact that vulnerability-prone patches correlate with a small set of observable code and task-related characteristics suggests a path toward lightweight, automated risk assessment for APR patches, analogous to prior work in abnormal commit detection (rosen2015commit; goyal2018identifying; leite2015uedashboard; gonzalez2021anomalicious). Simple signals such as how widely a patch is spread across the codebase and the type of available information in an issue description can be combined into a prediction model that scores risk of APR usage before and after patch generation. In the input, issue-level signals can be used to flag high-risk tasks for APR systems. In the output, an additional set of code-level features can be used to evaluate generated patches. A risk model leveraging these factors would not replace static analysis or code review, but can be integrated into CI/CD pipelines as an early warning and triage mechanism where high-risk APR tasks or patches get routed to stricter security checks or human review. The signals needed for such risk assessment are already present in typical development workflows, making the integration of risk assessment into CI/CD a promising direction for future work.
6. Related Work
Security Analyses of LLM-based APR Systems. Recent work has begun to analyze the security of LLM-based APR agents, rather than only their repair capabilities. Przymus et al. show that carefully crafted adversarial bug reports can mislead APR pipelines, inducing CVE reversions, vulnerability injection, CI/CD abuse, and denial-of-service even when some defenses such as AI-assisted reviews are deployed (przymus2025adversarialbugreportssecurity). Their results highlight the bug-report interface as a new attack surface and suggest that current defenses block only a subset of malicious reports.
Chen et al. conduct a complementary red-teaming study of APR agents in realistic CI/CD settings, asking whether seemingly valid issue statements can cause agents to generate vulnerable patches (chen2025redteamingprogramrepair). They challenge the common assumption that tests passing implies safety for APR-generated patches, and demonstrate that agents can be compelled into introducing exploitable behavior while still satisfying existing test suites.
Both works build on a rapidly growing ecosystem of LLM-based APR agents which primarily focus on improving repair success and benchmark performance in repository-scale settings, often without any security checks (ruiz2025art; yang2025enhancing; zhao2024enhancing; fan2023automated; zhang2024autocoderover; openhands; honeycomb). In contrast to these adversarial perspectives, our study analyzes LLM and agentic APR outputs under non-adversarial SWE-bench issues, quantifying how often vulnerabilities arise by accident and linking them to code and task-level factors such as edit scope and issue information completeness, as well as comparing them directly to developer-written patches.
@@ 6. Related Work
Security Analyses of LLM-generated Code. Given the growing adoption of LLMs by developers (daigle2024aiwave; shani2023aiimpact; kabir2024stack), recent studies have begun to examine various aspects of LLM-generated code (fan2023large; liu2024exploring; jimenez2024swebench; wang2025solved). While many of these studies have focused on functionality, others have examined the security implications of LLM-generated code (pearce2025asleep; fu2023security; toth2024llms; khoury2023secure; siddiq2024quality; mohsin2024can). Pearce et al. (pearce2025asleep) conducted one of the first assessments on the security of LLM-generated code, showing that even state-of-the-art models can produce vulnerable code in security-sensitive scenarios. Since then, many studies have confirmed these findings by analyzing LLM-generated code across different tasks, models, and user interaction settings (ramirez2024state; perry2023users; liu2024refining; siddiq2023generate).
@@
Most prior work has focused on small code snippets or controlled experimental settings (mohsin2024can; siddiq2024quality; pearce2025asleep; khoury2023secure). These settings often lack the full project context and complexity inherent in real-world software maintenance tasks, limiting their ability to assess how vulnerabilities emerge in practice. However, a recent study by Fu et al. (fu2023security) analyzed Copilot-generated code snippets in GitHub projects and identified several security weaknesses. It is worth mentioning that because these completions are reviewed and integrated by developers, the model is not fully in control of the final output. Our work complements and extends this work by analyzing LLM-generated code in real-world GitHub issues, where models are given autonomy to modify any relevant files. To our knowledge, this is the first large-scale empirical study to link vulnerabilities introduced by LLMs and agentic frameworks to a broad set of contextual factors, including issue descriptions, code-level attributes, and project structure, in realistic software development settings where models are tasked with resolving issues.
Vulnerability Detection. Automated vulnerability detection can be performed using static or dynamic analysis, with static analysis being widely adopted for large-scale, code-centric studies. static analysis detects weaknesses by reasoning about source code without execution through techniques such as rule/pattern matching, taint and dataflow analysis, and abstract interpretation (owasp_source_code_analysis). Tools like CodeQL, Semgrep, and SpotBugs leverage these techniques with CWE-based queries and rules that support automated evaluations of code (GitHubCodeQL; semgrep; spotbugs). More recent work has explored leveraging LLMs for detection, either directly, by identifying vulnerabilities in code snippets or projects (ullah2024llms; chakraborty2021deep), or indirectly by enhancing static analysis pipelines (sheng2025llms; lu2024grace). In the latter case, LLMs have been used to tasks such as infering vulnerability-specific taint specifications, identifying function call chains, and reasoning over dataflow paths to improve static taint analysis (li2025iris; li2025hitchhiker; liu2025llm). While LLM-based detection introduces new opportunities, established static analysis tools remain widely used in practice, particularly in CI/CD pipelines, where their deterministic query semantics and reproducibility are often essential.
@@
Abnormal Commit Detection. Past studies have explored various techniques to identify and characterize the risk of potential changes in software systems, particularly focusing on GitHub commits (rosen2015commit; goyal2018identifying; leite2015uedashboard; gonzalez2021anomalicious). This line of research will generally aim to identify outlier commits, also known as risky or abnormal commits, through assessing various commit attributes. For instance, Gonzalez et al. introduced Anomalicious (gonzalez2021anomalicious), an approach to identify malicious commits. By analyzing the commits and identifying the characteristics of a “normal” developer commit, they made an attempt at detecting malicious code contributions.
@@
7. Conclusions
This study presents the first large-scale security evaluation of an LLM and three agent-generated patches for real-world issue resolution tasks. We find that Llama can introduce up to 11x more new vulnerabilities than developers when addressing issues. Notably, some of the most frequent vulnerabilities introduced by Llama (e.g., CWE-95, CWE-502) differ significantly from those found in developer-written code (e.g., CWE-732, CWE-377), indicating that the model is not merely mimicking human mistakes but also introducing distinct patterns of failure. Further, our analysis of agentic frameworks shows that increasing LLM autonomy can further amplify vulnerability risks, especially in tasks that involve development of new features or tasks that require cross-file reasoning.
Importantly, our findings show that vulnerable outputs are not random. Rather, they are associated with particular code and issue-level characteristics, such as a higher number of modified files, number of unique file types, and the absence of code snippets or certain information in the issue description. These patterns suggest that the riskiness of LLM or agent-generated patches can be assessed not only by detecting specific vulnerable lines of code but also by developing complementary risk assessment techniques that evaluate both the generated code and the broader context in which a patch is proposed e.g., issue type or information completeness. Such systems can complement existing vulnerability detection techniques, often reliant on costly manual reviews, by narrowing the focus to potentially high-risk code contributions.